OpenStudy (anonymous):

what is the slope of a line that passes through the points (1,-5) and (-3,3)?

OpenStudy (tkhunny):

Have you considered calculating it? \(\dfrac{3 - (-5)}{-3 - (1)}\) -- Go!

OpenStudy (nincompoop):

IF you were given two points (a, b) and (c, d) first, find the slope of the line that connects them slope is the "steepness" of a line in layman's terms, but more formally the ratio of change in y and x \[slope = m = \frac{\Delta y}{\Delta x}=\frac{y_{final}-y_{initial}}{x_{final}-x_{initial}}=\frac{d-b}{c-a}\]

OpenStudy (anonymous):